# TextMine: LLM-Powered Knowledge Extraction for Humanitarian Mine Action

### v1

@@ 4 Experimental Results
4.1 Test Dataset
We constructed a dataset based on six recent mine action reports from CMAC: the 2023 Annual Report, the 2023–2024 Integrated Work Plans, the 2023 Cluster Munition Remnants Clearing Report, the 2023 Mine Clearing Report, and the Article 7 Report. In total, we processed 270 pages, yielding 549 text chunks. These text chunks are used as test instances in zero-shot prompts. We then generate model responses using GPT-4o and Llama3-70B for these zero-shot prompts. 100 prompt-response pairs are randomly sampled from them. The sampling ensures a balanced distribution across ontology categories.
The annotation workflow involved a human annotator reviewing the model-generated triples and discarding those that were semantically or factually incorrect. To prevent any bias, the annotator was kept blind to which model produced each output, ensuring that validation decisions did not favor a particular system. The remaining valid triples were then aggregated and de-duplicated to ensure uniqueness. This process resulted in a final dataset of 1,095 unique triples derived from 100 sampled prompt-response pairs. This curated dataset serves as a reliable test set for model evaluation by minimizing noise and redundancy while maintaining ontology diversity.
In the retrieval-based one-shot prompt template construction phase, some examples were retrieved from this dataset to construct demonstrations. These retrieved examples were excluded from being sampled into the final dataset to ensure an unbiased evaluation.
4.2 Reference-Based Evaluation
@@
Figure 4: The combined visualization illustrates the impact of model selection and prompt strategy on extraction performance. The top Combined Score is achieved by Llama3-70B (93.24) closely followed by GPT-4o (93.13), both with OS prompt setting.
Figure 5 further breaks down the performance of five models across four accuracy evaluation metrics: BLEU, ROUGE, METEOR, and BERTScore. OS demonstration prompts consistently result in the best accuracy across all four metrics and all five models, highlighting their effectiveness for the triple extraction task. BLEU scores peak with OS prompts, with GPT-4o achieving the best performance. ROUGE results show a similar trend, with GPT-4o and Llama3-70B excelling in the OS prompt setting. METEOR follows the same pattern as BLEU and ROUGE, reinforcing the advantages of OS prompts. BERTScore, which measures semantic similarity, shows high clustering across models, suggesting minimal differentiation in performance. Overall, OS demonstration prompts consistently enhance extraction accuracy across all models and metrics.
@@
Figure 6 presents the hallucination rates for subjects, relations, and objects across different models and prompt settings. GPT-4o, Llama3-70B, and Gemma2-9B exhibit lower hallucination rates for subjects and objects, while Llama3-8B and Mistral-7B tend to have higher hallucination rates. For high-performing models, the OS prompt type generally helps reduce hallucination, whereas RP prompts tend to increase it. Zero-Shot prompts often lead to increased hallucination for subjects and objects across models. Interestingly, however, Zero-Shot prompts show lower hallucination rates for relations, which may be due to the additional demonstrations in other prompts introducing noise that negatively impacts relation extraction.
@@
Most models exhibit high format conformance across all prompt types, with only Gemma-9B under the Zero-Shot prompt scoring below 80%. GPT-4o and Llama3-70B consistently achieve FC above 95% across all prompt types, demonstrating superior adherence to the expected format.
4.3 Reference-Free Evaluation
Reference-based evaluation shows that the OS prompt consistently achieves the highest performance across most models, so for the reference-free ranking experiments we only focus on ranking the five models under OS prompt setting. To assess the alignment between reference-based and reference-free rankings, we compute correlations between the rankings derived from the Combined Score calculated based on references, and the rankings derived from the Expectation Score based on LLM judges.
@@ 4.3 Reference-Free Evaluation
Table 3: Correlation values for different judge models and judge methods.
@@ 5 Related Work
The emergence of LLMs such as those from the GPT family (e.g., GPT-4o [^11] OpenAI GPT-4o, https://openai.com/index/hello-gpt-4o/), Llama (Touvron et al., 2023), BLOOM (Workshop et al., 2022), and PaLM (Chowdhery et al., 2023), has significantly transformed the field of knowledge extraction from text. These models possess advanced language understanding and reasoning capabilities, making them well-suited for extracting knowledge from unstructured text or well-structured documents Colakoglu et al. (2025), especially when paired with prompting techniques like in-context learning (ICL) (Brown et al., 2020).
ICL enables LLMs to learn new tasks by providing input-output demonstrations during inference. Depending on the number of examples provided, this can range from zero-shot (no demonstrations) to one-shot or few-shot learning (multiple demonstrations) (Min et al., 2022; Liu et al., 2023). This method enhances the ability of the models to generalize from minimal data. Zhu et al. (2023) demonstrated the effectiveness of ICL through a virtual knowledge extraction task.
Mihindukulasooriya et al. (2023) introduced an approach that uses ontology guidance to extract knowledge from text. Their work highlights the potential of LLMs in extracting domain-specific knowledge constrained by ontological rules. In our study, we adapt this approach to the domain of humanitarian demining, employing a set of specialized ontologies to guide the extraction process.
Evaluating generated texts is a challenging task, especially when limited ground-truth data are available. To address this problem, recent approaches include generating synthetic data to train an evaluator model (Saad-Falcon et al., 2023), annotating datasets using a human-in-the-loop methodology (Dagdelen et al., 2024), or leveraging strong LLMs as judges (Zheng et al., 2024; Bavaresco et al., 2024). Our work involves annotating extracted triples to create an evaluation dataset, applying LLMs as judges and analyzing the alignment between these two evaluation methods.
@@ 6 Conclusion
TextMine addresses the need for automated knowledge extraction from HMA reports by LLMs and domain ontologies to transform unstructured reports into structured knowledge triples. Our findings highlight the effectiveness of ontology-aligned prompts in improving extraction accuracy. The introduction of the HMA ontology and a human-annotated evaluation dataset bridges the gap in standardized resources for the HMA domain. Furthermore, our bias-aware reference-free evaluation using LLM-as-a-Judge demonstrates great potential.
@@
While TextMine demonstrates the feasibility of an LLM-driven extraction pipeline in the specialized domain of humanitarian demining, several limitations warrant acknowledgement. First, our evaluation relies on just 100 prompt–response examples (yielding 1,095 unique triples); although these examples were selected to cover diverse geographies and operational contexts, the small scale constrains broad generalizability. Assembling and annotating demining data demands extensive domain expertise, so even this modest set offers valuable proof-of-concept insights, but we plan to extend to larger, multilingual collections in future work. Second, ground-truth triples were produced via a model-assisted, expert-validated workflow by a single annotator; the absence of multiple annotators prevents calculation of inter-annotator agreement, and future studies will involve multi-annotator labeling to quantify label reliability. Third, although the annotator was blinded to model provenance when validating candidate triples, using model suggestions as a starting point may introduce subtle bias toward the evaluated model family; to mitigate this, we intend to curate a purely human-authored gold set for a representative subset. Finally, while we acknowledge that in-depth analysis of the LLM-judge’s own hallucination and bias characteristics is important, conducting a comprehensive bias and hallucination audit falls beyond this paper’s core scope and is deferred to future work focused explicitly on judge calibration and fairness.
@@
Appendix A In-Context Learning Prompt Example
@@
The below is an example of one-shot prompt with RS prompt setting. All in-context learning prompts are stored in CSV files as part of the supplementary material for easier reproducibility.
@@
Appendix B LLM Judge Prompts
@@
Appendix C Evaluation Reasoning Process

### v2

@@ 4 Experimental Results
Reference-Based Evaluation
@@
Figure 4: The combined visualization illustrates the impact of model selection and prompt strategy on extraction performance. Abbriviations on x-axis are about One-shot with: RS = Random Sentences; RP = Random Paragraphs; OS = Ontology-Aligned Sentence; OP = Ontology-Aligned Paragraph. The top Combined Score is achieved by Llama3-70B (93.24) closely followed by GPT-4o (93.13), both with OS prompt setting.
Figure 5 further breaks down the performance of five models across four accuracy evaluation metrics: BLEU, ROUGE, METEOR, and BERTScore. OS demonstration prompts consistently result in the best accuracy across all four metrics and all five models, highlighting their effectiveness for the triple extraction task. BLEU scores peak with OS prompts, with GPT-4o achieving the best performance. ROUGE results show a similar trend, with GPT-4o and Llama3-70B excelling in the OS prompt setting. METEOR follows the same pattern as BLEU and ROUGE, reinforcing the advantages of OS prompts. BERTScore, which measures semantic similarity, shows high clustering across models, suggesting minimal differentiation in performance.
Overall, OS demonstration prompts consistently enhance extraction accuracy across all models and metrics.
@@
Figure 6 presents the hallucination rates for subjects, relations, and objects across different models and prompt settings. GPT-4o, Llama3-70B, and Gemma2-9B exhibit lower hallucination rates for subjects and objects, while Llama3-8B and Mistral-7B tend to have higher hallucination rates. For high-performing models, the OS prompt type generally helps reduce hallucination, whereas RP prompts tend to increase it. Zero-Shot prompts often lead to increased hallucination for subjects and objects across models. Interestingly, however, Zero-Shot prompts show lower hallucination rates for relations, which may be due to the additional demonstrations in other prompts introducing noise that negatively impacts relation extraction. Most models exhibit high format conformance across all prompt types, with only Gemma-9B under the Zero-Shot prompt scoring below 80%. GPT-4o and Llama3-70B consistently achieve FC above 95% across all prompt types, demonstrating superior adherence to the expected format.
@@
Reference-Free Evaluation Results show that the OS prompt consistently achieves the highest performance across most models, so for the reference-free ranking experiments we only consider five models under OS prompt setting. To assess the alignment between reference-based and reference-free rankings, we compute correlations between the rankings derived from the Combined Score calculated based on references, and the rankings derived from the Expectation Score based on LLM judges.
@@
Table 2: The correlation results for different judge models, judge methods.
@@ 5 Related Work
LLMs, such as those from the GPT family Hurst et al. (2024), Llama (Touvron et al., 2023), BLOOM (Workshop et al., 2022), and PaLM (Chowdhery et al., 2023), have transformed the field of knowledge extraction from text. They possess advanced language understanding and reasoning capabilities, making them well-suited for extracting knowledge from unstructured text or well-structured documents Colakoglu et al. (2025), especially when paired with prompting techniques like in-context learning (ICL) (Brown et al., 2020).
ICL enables LLMs to learn new tasks by providing input-output demonstrations during inference. Depending on the number of examples provided, this can range from zero-shot (no demonstrations) to one-shot or few-shot learning (multiple demonstrations) (Min et al., 2022; Liu et al., 2023b). This method enhances the models’ ability to generalize from minimal data. Zhu et al. (2023) showed the effectiveness of ICL for knowledge extraction. Mihindukulasooriya et al. (2023) introduced an approach that utilizes ontology guidance to extract knowledge from text. Their work highlights the potential of LLMs in extracting domain-specific knowledge constrained by ontological rules. In our study, we adapt this approach to the domain of humanitarian demining, employing a set of specialized ontologies to guide the extraction process.
Evaluating generated texts is a challenging task, especially when limited ground-truth data are available. To address this problem, recent approaches include multi-faceted fact-based evaluation Gashteovski et al. (2022), generating synthetic data to train an evaluator model (Saad-Falcon et al., 2023; Kim et al., 2025), annotating datasets using a human-in-the-loop methodology (Dagdelen et al., 2024), or leveraging strong LLMs as judges (Zheng et al., 2024; Bavaresco et al., 2024). Our work involves annotating extracted triples to create an evaluation dataset, applying LLMs as judges and analyzing the alignment between these two evaluation methods. For more detailed discussion on related work, see Appendix A.
@@ 6 Conclusion
TextMine addresses the need for knowledge extraction from HMA reports by LLMs and domain ontologies to transform unstructured technical reports into structured knowledge triples. This paper introduces the curated data of HMA reports, ontology, and LLM pipeline, aiming to enable innovations for mine action and natural language processing.
@@
While TextMine provides the first dataset, evaluation framework and pipeline that demonstrates the feasibility of an LLM-driven extraction pipeline in the specialized domain of humanitarian demining, several limitations are considered to be addressed in the future, including the further extension of the dataset. Our evaluation relies on 100 prompt–response examples (yielding 1,095 unique triples). Assembling and annotating demining data demands extensive domain expertise, so even this modest set offers valuable proof-of-concept insights, but we plan to extend to larger, multilingual collections in future work. Second, this work is focues on reports that are written only on English. In the future, we plan to extend the work to include other languages, because mine-action data are often produced in multiple languages, such as Arabic, French or Ukrainian, depending on the conflict region. Making TextMine multilingual is, therefore, essential to ensure global applicability and inclusiveness.
@@
Appendix A Detailed Discussion of Related Work
A.1 Automated Knowledge Graph Construction
Constructing knowledge graphs (KGs) from text is a well-established problem Weikum et al. (2021). The methods for constructing KGs are either schema-free Pei et al. (2023); Kotnis et al. (2022b); Gashteovski et al. (2019) or are fixed either to general-purpose schemas Zhang and Soh (2024); Josifoski et al. (2022); Broscheit et al. (2017) or to domain-specific schemas Dagdelen et al. (2024); Lauscher et al. (2019); Broscheit et al. (2017). Both the schema-free methods and the methods that are tied to general-purpose schemas do not address the complexities of the domain at hand, which is why we worked on a specific data, benchmarks and methods for the HMA domain.
A.2 LLMs for Knowledge Graph Construction
Prior work has explored the use of LLMs for constructing KGs from text Pan et al. (2023). One line of work focuses on harvesting knowledge from pretrained LLMs without the use of annotated data via zero-shot iterative prompting Hu et al. (2025); Carta et al. (2023); Hao et al. (2022) or in-context learning (ICL) Brown et al. (2020); Min et al. (2022); Liu et al. (2023a). However, these methods are not well-suited for the construction of domain-specific knowledge graphs.
For these reasons, there have been methods for constructing domain-specific KGs by the use of LLMs Chen et al. (2024). For example, there are LLM methods for constructing KGs in various domains, such as medicine Arsenyan et al. (2023), finance Li et al. (2025), cyber-thread intelligence Huang and Xiao (2024) or climate science Pan et al. (2025). To the best of our knowledge, however, there is no prior work on KG construction for the humanitarian mine action domain.
A.3 Ontology-Guided for Knowledge Graph Construction
Ontology-guided (a.k.a. schema-driven) methods have been shown to be effective, though they are often limited to single-sentence input and, consequently, overlook context-level reasoning Cauter and Yakovets (2024). Text2KGBench Mihindukulasooriya et al. (2023) is a benchmark and evaluation framework for extracting triples from text, which are bound to a predefined ontology. However, it relies on toy data schemas, which limits their applicability to the complex, domain-specific HMA documents.
Other line of work focuses on reasoning and validation within already structured knowledge graphs, not extraction from free-form text as with TextMine Liu et al. (2025); Regino and Dos Reis (2025). Their ontological constraints apply post-hoc to existing triples, whereas TextMine uses the ontology during extraction to guide triple formation and reduce hallucinations.
Finally, none of these works address bias-aware evaluation: they rely on deterministic metrics or small-scale manual validation. TextMine introduces a bias-aware LLM-as-Judge evaluation framework that explicitly mitigates positional bias in reference-free scoring—an issue absent from prior ontology-guided pipelines.
A.4 Multi-Faceted Knowledge Graph Evaluation
They are typically evaluated with a single score metric, such as accuracy Eisenstein (2019). In recent years, however, the NLP community has increasingly been pointing out that such evaluation approach can obscure the model performance nuances, thus revealing misleading results Ribeiro et al. (2020).
To mitigate these effects, researchers have turned to multiple metric evaluations, which reveal more complete picture about the performance of the models Vickers et al. (2024); Opitz (2024); Jain et al. (2023); Deutsch et al. (2022); Peyrard et al. (2021); Liu et al. (2021). Particularly within the field of information extraction, such multi-faceted evaluations are already well established Radevski et al. (2023); Friedrich et al. (2022); Gashteovski et al. (2022); Fu et al. (2020). Likewise, KG-related work also employs such multi-faceted evaluations, such as for the task of link prediction Gastinger et al. (2023); Widjaja et al. (2022); Rim et al. (2021); Meilicke et al. (2018).
In order to evaluate aspects of the models that are important to users or that need immediate human verifications, another line of work have used manual evaluations Bojic et al. (2023); Novikova et al. (2018). This could be done with user studies that target the specific aspects that one tries to measure Radevski et al. (2025); Schuff et al. (2023); Kotnis et al. (2022a) or by simply manually validating the final results Thomson et al. (2024); van der Lee et al. (2021). Such manual evaluations have been practiced for both information extraction research works Sainz et al. (2023); Gashteovski et al. (2020); Bhardwaj et al. (2019), as well as for KG prediction tasks, such as link prediction Carriero et al. (2024); Xu et al. (2024a); Zhou et al. (2022). Manual evaluations, however, are both expensive and not scalable, thus making them practically infeasible.
To avoid the pitfalls of the narrow single-score metrics, as well as the scalability issues of manual evaluations, we propose a multi-faceted evaluation framework that incorporates several aspects of the information extraction problem, fused into the final score.
Appendix B Additional Information and Insights for the HMA Ontology
We visualize the statistics regarding the combined ontologies. Fig. 7 includes the number of prompts, entity types, and relation types per ontology.
The Empathi Gaur et al. (2019) ontology is designed for “Emergency Managing and Planning about Hazard Crises”. It is the most advanced ontology that is openly-available in the research domain for hazard crises, as an outcome of a research study. However, this ontology is designed for the broader concept of disasters as opposed to being specialized in the HMA domain. For instance, it has many concepts such as hurricanes or firefighters that are outside of the scope of the demining-relevant disasters such as mine explosions. The concepts (entity types and relation types) that are relevant to the demining are manually filtered and used as the ontology. Thus, the number of entities and relations are after the filtering process. GICHD’s new data schema IMSMA Core (previous standard was called IMSMA-NG Desantis and Eriksson (2013), where NG stands for “Next Generation”) is the agreed standard between mine agencies of governments and non-governmental organizations (NGOs). IMSMA Core serves the data model for databases of mine agencies globally. On the other hand, the number of entities and relation types are limited.
We consider the separate portions of the GICHD data model, which are all related to the mine action but logically separated as “Assistance, Accident, Activity, Land, Location, and Victim” ontologies. We utilize both as the best state-of-the-art models and convert them into relatively small ontologies. On the other hand, we utilize the concepts from all of the 7 ontologies in our prompt generation process for LLMs, resulting in different number of prompts per ontology. The proposed “HMA Ontology” can be extended with more special entity and relation types based on special needs of the organizations.
Figure 7: Number of prompts, entity types, and relation types per the utilized ontologies from IMSMA Core and Empathi.
Appendix C In-Context Learning Prompt Example
@@
The below is an example of one-shot prompt with the RS prompt setting. All in-context learning prompts are stored in CSV files as part of the supplementary material for easier reproducibility.
@@
Appendix D LLM Judge Prompts
@@
Appendix E Evaluation Reasoning Process